OpenStudy (anonymous):

Two similar pyramids have lateral areas 20ft^2 and 45^2 .The volume of the smaller pyramid is 8ft^3. What is the volume of the larger pyramid? A) 27ft^3 B) 18ft^3 C) 36ft^3 D) 9ft^3

OpenStudy (dumbcow):

ratio of areas is 20/45 = 4/9 ratio of side lengths is square root of area ratio: 2/3 ratio of volumes is ratio of sides cubed: = 2^3/3^3 = 8/27 thus volume of large pyramid is 27